OpenStudy (anonymous):

Determine the type of boundary line and shading for the graph of the inequality -5x - y (less than or equal to) -2 A) Dashed line with shading on the side that includes the origin. B) Solid line with shading on the side that does not include the origin. C) Dashed line with shading on the side that does not include the origin. D) Solid line with shading on the side that includes the origin. I know its either B or D ... /:

OpenStudy (anonymous):

good start, it has to be solid because we are including the value ( hence =) What is the origin in terms of a point?

OpenStudy (anonymous):

honestly I'm not sure, i just know that to get ready to graph it, it would be y (less than or equal to) -5x +2 ... i think

OpenStudy (anonymous):

the origin is the point (0,0) which is (x,y). plugging in 0 for x and y you get -5(0) -(0) <= -2 or 0<=-2 which is false, so the point is NOT included

OpenStudy (anonymous):

y < 2x + 4 This inequality would include it though right?

OpenStudy (maheshmeghwal9):

\[-5x-y \le -2.\]\[\implies 5x+y \ge 2.\] See the inequality Plot Just click the link:) http://www.wolframalpha.com/input/?i=-5x-y+%3C%3D-2

OpenStudy (anonymous):

@ CHarity, yes

OpenStudy (anonymous):

k just double checking haha thanks!

OpenStudy (maheshmeghwal9):

:D